OpenStudy (anonymous):

write ln as a sum and difference for logarithms ln(x^5y^3/z^2)

OpenStudy (zarkon):

\[\ln(x^5)+\ln(y^3)-\ln(z^2)\]

OpenStudy (zarkon):

\[5\ln(x)+3\ln(y)-2\ln(z)\]

OpenStudy (anonymous):

5ln(x)+3ln(y)-2ln(z)

OpenStudy (anonymous):

are both answers right?

OpenStudy (anonymous):

how did u arrive at that conclusion?

OpenStudy (anonymous):

both answer is correct

OpenStudy (anonymous):

Thank u!! GOD bless<><